# TRACE: Temporal Grounding Video LLM via Causal Event Modeling

### v1

@@ 2 Related Works
Large language models (LLMs) (Kaplan et al., 2020; Achiam et al., 2023; Touvron et al., 2023) have demonstrated significant potential in acquiring knowledge and addressing real-world challenges using a zero-shot approach. Recent research has focused on integrating knowledge from other modalities, such as vision (Liu et al., 2024; Li et al., 2023a) and audio (Ghosal et al., 2023), to bolster the capabilities of LLMs. Within the visual domain, video large language models (video LLMs) have emerged as a crucial research area (Lin et al., 2023a; Maaz et al., 2023; Zhu et al., 2023). Traditional video LLMs (Zhang et al., 2023; Lin et al., 2023a; Li et al., 2023b, 2024; Cheng et al., 2024; Yao et al., 2024) have made considerable performance improvements in tasks such as video question answering, reasoning, and video captioning. However, these methods encounter difficulties in precisely pinpointing event timestamps within videos. To address this issue, TimeChat (Ren et al., 2023), VTimeLLM (Huang et al., 2023), and Hawkeye (Wang et al., 2024b) have attempted to overcome this limitation by fine-tuning the video LLMs on VTG datasets. More recently, LITA (Huang et al., 2024) introduces fast-slow visual tokens and incorporates time tokens into LLM tokenizers. Momentor (Qian et al., 2024) suggests a time encoder to address time token quantization errors. VTG-LLM (Guo et al., 2024) integrates special time tokens and time position embeddings to improve the ability of video LLMs in comprehending timestamps. However, these methods do not take into account the inherent structure of videos and still cannot achieve satisfactory performance. In this paper, we propose the causal event modeling framework to model video structure and design the TRACE model to address the proposed framework. Numerical results demonstrate significant performance gains of TRACE over existing video LLMs on VTG tasks.
@@ 3.3 Training Strategy and Data Preparation
In stage 1, task modules such as the vision compression layer, time encoder/head, score encoder/head, and text tokenizer/head are trained while the vision encoder and LLM backbone remain fixed. As shown in Table 1, stage 1 primarily utilizes two groups of datasets.
  • Image and video caption datasets for initializing the visual compression layer. This group of datasets including Valley (Luo et al., 2023b), LLaVA_Image (Liu et al., 2024), TextVR (Wu et al., 2025), and a randomly sampled subset of ShareGPT4Video (Chen et al., 2024a) datasets.
@@
  • Video question answering datasets to enhance TRACE’s reasoning capabilities. We use VideoChatGPT (Maaz et al., 2023) and Next-QA (Xiao et al., 2021) in this part.
@@
Table 1: Datasets used for TRACE training process. "Compressed" indicates that datasets are condensed by retaining only one sample for samples with identical videos but varying instructions.
@@
Stage 2 Valley (Compressed), TextVR (Compressed), ShareGPT4Video (Compressed), VTG-IT, ActivityNet Captions, VideoChatGPT, InternVid, Next-QA 0.9M
@@ 4.2 Performance of TRACE
In Table 4, we show the performance of TRACE on ActivityNet Captions dataset. All the reported algorithms except for Momentor (Qian et al., 2024) have incorporated the ActivityNet Captions dataset as part of the training data. Results show that the TRACE attains the best performance in moment retrieval tasks and demonstrates comparable results to VTimeLLM in dense video caption tasks.
@@ A.1 Details of data format
  • General tasks (Figure 6). For general tasks such as video captioning, image captioning, and video question answering, the answer component of the data does not include timestamps or scores. Consequently, we employ a single token $\langle sync\rangle$ as a placeholder for timestamps and scores, signifying an empty response for this part of response. This kind of annotation including LLaVA_Image (Liu et al., 2024), Valley (Luo et al., 2023b), TextVR (Wu et al., 2025), ShareGPT4Video (Chen et al., 2024a), VideoChatGPT (Maaz et al., 2023), and Next-QA (Xiao et al., 2021) datasets.
@@ B.2 Additional Experiment Results
In Tables 7 and  8, we report the performance of TRACE using different training data. The results show that (1) the performance of TRACE on long videos increase while using the original TRACE setting. (2) The performance of TRACE slightly reduced on short videos (Charades-STA).

### v2

@@ 2 Related Works
Large language models (LLMs) (Kaplan et al., 2020; Achiam et al., 2023; Touvron et al., 2023) have demonstrated significant potential in acquiring knowledge and addressing real-world challenges using a zero-shot approach. Recent research has focused on integrating knowledge from other modalities, such as vision (Liu et al., 2024b; Li et al., 2023a) and audio (Ghosal et al., 2023), to bolster the capabilities of LLMs. Within the visual domain, video large language models (video LLMs) have emerged as a crucial research area (Lin et al., 2023a; Maaz et al., 2023; Zhu et al., 2023). Traditional video LLMs (Zhang et al., 2023; Lin et al., 2023a; Li et al., 2023b, 2024a; Cheng et al., 2024; Yao et al., 2024) have made considerable performance improvements in tasks such as video question answering, reasoning, and video captioning. However, these methods encounter difficulties in precisely pinpointing event timestamps within videos. To address this issue, TimeChat (Ren et al., 2023), VTimeLLM (Huang et al., 2023), and Hawkeye (Wang et al., 2024b) have attempted to overcome this limitation by fine-tuning the video LLMs on VTG datasets. More recently, LITA (Huang et al., 2024) introduces fast-slow visual tokens and incorporates time tokens into LLM tokenizers. Momentor (Qian et al., 2024) suggests a time encoder to address time token quantization errors. VTG-LLM (Guo et al., 2024) integrates special time tokens and time position embeddings to improve the ability of video LLMs in comprehending timestamps. However, these methods do not take into account the inherent structure of videos and still cannot achieve satisfactory performance. In this paper, we propose the causal event modeling framework to model video structure and design the TRACE model to address the proposed framework. Numerical results demonstrate significant performance gains of TRACE over existing video LLMs on VTG tasks.
@@ 3.3 Training Strategy and Data Preparation
In stage 1, task modules such as the vision compression layer, time encoder/head, and score encoder/head are trained while the vision encoder and LLM backbone remain fixed. As shown in Table 1, stage 1 primarily utilizes two groups of datasets.
  • Image and video caption datasets for initializing the visual compression layer. This group of datasets including Valley (Luo et al., 2023b), LLaVA_Image (Liu et al., 2024b), TextVR (Wu et al., 2025), and a randomly sampled subset of ShareGPT4Video (Chen et al., 2024a) datasets.
@@
  • Video question answering datasets to enhance TRACE’s reasoning capabilities. We use VideoChatGPT (Maaz et al., 2023) and Next-QA (Xiao et al., 2021) in this part. For TRACE-uni, we have incorporated the YouTube and perception-test components from LLaVA-Video-178K (Zhang et al., 2024).
@@
Table 1: Datasets used for TRACE training process. "Compressed" indicates that datasets are condensed by retaining only one sample for samples with identical videos but varying instructions. The transparent text denotes the setting of TRACE-uni, which involves the inclusion of additional training data pertaining to general video understanding tasks.
@@
Stage 2 Valley (Compressed), TextVR (Compressed), ShareGPT4Video (Compressed), VTG-IT, ActivityNet Captions, VideoChatGPT, InternVid, Next-QA, LLaVA-Video-178K 0.9M (1.77M)
@@ 4.2 Performance of TRACE
  • Effectiveness of the TRACE architecture in addressing general video understanding tasks. As illustrated in Table 6 and 2, we observe that: (1) despite not being trained on extensive multi-task datasets, TRACE is still effective in managing general QA tasks; (2) by incorporating additional training data for general video understanding tasks, the TRACE-uni achieves superior performance compared to TRACE in both VTG tasks and general video understanding tasks.
@@
In Table 4, we show the performance of TRACE on ActivityNet Captions dataset. All the reported algorithms except for Momentor (Qian et al., 2024) have incorporated the ActivityNet Captions dataset as part of the training data. Results show that the TRACE attains the best performance in both moment retrieval tasks and dense video caption tasks.
@@
Table 6: Performance of TRACE on general video understanding tasks.
Models MVBench (Li et al., 2024b) VideoMME (w/o subtitle) (Fu et al., 2024)
Video-LLaVA (Lin et al., 2023a) 41.0 39.9
LLaVA-NeXT-Video (Liu et al., 2024a) 46.5 33.7
VideoChat2 (Li et al., 2024b) 51.1 33.7
VideoLLaMA 2 (Cheng et al., 2024) 54.6 44.6
TRACE 48.1 43.8
TRACE-uni 53.8 49.6
@@ A.1 Details of data format
  • General tasks (Figure 6). For general tasks such as video captioning, image captioning, and video question answering, the answer component of the data does not include timestamps or scores. Consequently, we employ a single token $\langle sync\rangle$ as a placeholder for timestamps and scores, signifying an empty response for this part of response. This kind of annotation including LLaVA_Image (Liu et al., 2024b), Valley (Luo et al., 2023b), TextVR (Wu et al., 2025), ShareGPT4Video (Chen et al., 2024a), VideoChatGPT (Maaz et al., 2023), and Next-QA (Xiao et al., 2021) datasets.
@@ B.2 Additional Experiment Results
In Tables 8 and  9, we report the performance of TRACE using different training data. The results show that (1) the performance of TRACE on long videos increase while using the original TRACE setting. (2) The performance of TRACE slightly reduced on short videos (Charades-STA).